OpenStudy (anonymous):

"One alloy contains 80 percent gold and another alloy contains 55 percent gold. How many grams of each alloy should be combined to make 40g of an alloy that contains 70 percent gold?"

OpenStudy (anonymous):

put the 80% alloy as x, then it has 80% of x = 0.8x ounces of gold the rest must be 40 - x and it has 55% gold so it has \[0.5(40-x)\] ounces gold. the total is 70% of 40 = .7 times 40 = 28 grams of gold solve \[0.8x+0.5(40-x)=28\] for x to get your answer

OpenStudy (anonymous):

that is wrong, it is 55% gold so solve \[0.8x+0.55(40-x)=28\]

OpenStudy (anonymous):

can you solve this or do you need help with it?

OpenStudy (anonymous):

need help please

OpenStudy (anonymous):

how to get the x?

OpenStudy (anonymous):

ok if you hate the decimals multiply both sides by 100 to get \[80x+55(40-x)=2800\] then multiply out to get \[80x+2200-55x=2800\] combine like terms to get \[25x+2200=2800\] subtract 2200 to get \[25x=600\] divide by 25 to get \[x=24\]

OpenStudy (anonymous):

thanks satellite

OpenStudy (anonymous):

yw